OpenStudy (anonymous):

Please help me with this algebra 1 problem! I will give a medal and follow. Solve: 8/x+15/x^2=-1 a. x=5,-3 b. x=-5,-3 c. x= -5, 3 d. x= 5,3 I know how to do this problem when numbers are the denominators with the xs. It's just the xs with no numbers and the one that is cubed really confuses me.

OpenStudy (anonymous):

@rishavraj

OpenStudy (anonymous):

*I meant squared

rishavraj (rishavraj):

\[\frac{ 8 }{ x } + \frac{ 15 }{ x^2 } = -1\] so get the LCM of denominators i.e x^2 so it would be \[\frac{ 8x + 15 }{ x^2 } = -1 \]

OpenStudy (anonymous):

so then do i cross multiply?

OpenStudy (anonymous):

what happened to the other x?

rishavraj (rishavraj):

yeah now u need to cross multiply...... for the another x.....hmmmmmm u tell me how u slove 8/3 + 7/9